The Red Crescent Society Association reports that the average time it takes to respond to an emergency call is 25 minutes. Assume the variable is approximately normally distributed and the standard deviation is 4.5 minutes. Calculate the probability that the time it takes to respond to an emergency call in less than 15 minutes.



1
Expert's answer
2021-07-05T14:43:41-0400

"\mu=25 \\
\sigma=4.5 \\
P(X<15) = P(Z< \frac{15-25}{4.5}) \\
= P(Z< -2.22) \\
= 0.013"